OpenStudy (anonymous):

A parachutist's rate during a free fall reaches miles per hour. What is this rate in feet per second? At this rate, how many feet will the parachutist fall during seconds of free fall? In your computations, use the fact that mile is equal to feet. Do not round your answers.

OpenStudy (anonymous):

To find the rate in feet/second, you need to multiply the speed in miles per hour by 30/44. This is a simplification of\[x \frac{miles}{hour}*\frac{5280~ft}{1~mile}*\frac{1~hour}{3600~\sec}=y \frac{ft}{\sec}\]

OpenStudy (anonymous):

Correction to above: 44/30.